OpenStudy (anonymous):

convert the following complex number into its polar representation 2+2i

OpenStudy (anonymous):

\[a+bi=r\left(\cos(\theta)+i\sin(\theta)\right)\] what you need is \(r\) and \(\theta\) \[r=\sqrt{a^2+b^2}\]

OpenStudy (anonymous):

and in this case (although not always), \(\theta=\tan^{-1}\left(\frac{b}{a}\right)\)

OpenStudy (anonymous):

2\[2\sqrt{2}\]? then what

OpenStudy (anonymous):

find \(\theta\) via \(\theta=\tan^{-1}\left(\frac{b}{a}\right)\)

OpenStudy (anonymous):

or you can just find \(\theta\) in this case by thinking what the angle has to be 

OpenStudy (anonymous):

i got .7853 for tan

OpenStudy (anonymous):

i would go for \(\frac{\pi}{4}\)

OpenStudy (anonymous):

okay then what after that

OpenStudy (anonymous):

then you are done

OpenStudy (anonymous):

thank you i se

OpenStudy (anonymous):

see

OpenStudy (anonymous):

don't forget this equality is between numbers, one number equals another number

OpenStudy (anonymous):

\[2+2i=2\sqrt2\left(\cos(\frac{\pi}{4})+i\sin(\frac{\pi}{4})\right)\]

OpenStudy (anonymous):

\[2\sqrt2\left(\cos(\frac{\pi}{4})+i\sin(\frac{\pi}{4})\right)\] \[=2\sqrt2(\frac{\sqrt2}{2}+i\frac{\sqrt2}{2})\] \[=2+2i\]